OpenStudy (anonymous):

The combined area of Jasper National Park and Banff National Park is about 17 500 km. The area of Banff NAtional park is 900 km greater than half the area of Jasper National Park. Find the area of Banff national Park.

OpenStudy (anonymous):

Let x be equal to the area of Jasper, and y be equal to the area of Banff. We then have: x+y=17500 y-x=900 We can rearrange equation 2 to be y=900+x Substitute into equation 1 and we get x+900+x=17500 This becomes 2x=16600 So x=8300 km Then y=8300+900=9200 km So the area of Banff is 9200 km

OpenStudy (anonymous):

But what about the "The area of Banff NAtional park is 900 km greater than HALF THE AREA of Jasper National Park" sorry for caps idk how to underline

OpenStudy (anonymous):

Lol btw I was born in Markham :P nice to know your from Mississauga

OpenStudy (anonymous):

Oh shoot, didn't see the half. Allow me to fix that: x+y=17500 y-x/2=900 Proceed as in my initial reply, and you will eventually get: Something to the effect of 1.5x=16600 So area of Banff is approximately 6433.3 km... This is closer to the actual value according to Wikipedia.

OpenStudy (anonymous):

Thanks again!

OpenStudy (anonymous):

No problem!